OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x - 3)(x + 1)(x - 5)?

OpenStudy (anonymous):

This function is already factored, so we don't have to take that step. To find the zeroes (roots) we set the function equal to 0, so: (x-3)(x+1)(x-5)=0 In order for this function to equal 0, one of the factors must be equal to 0 (you can't multiply two numbers together to get 0 if neither one is 0). So you can set each factor equal to 0: x-3=0 x+1=0 x-5=0 Solve each of these three expressions to get your roots.

OpenStudy (anonymous):

x=3 x=-1 x=5

OpenStudy (anonymous):

Its always the easy one that get me. Thanks!

OpenStudy (anonymous):

You're welcome, glad I could help